### v1

@@ 0.4.1 Generalisability-Verifiability Trade-off
To give us a sense of progression from the largest hyper-rectangle to the smallest, we also cluster $\mathbb{H}_{sh}$, see the last graph of Figure 3. For the clustering method, we employ the k-means algorithm to generate multiple ensembles of hyper-rectangles around each cluster. This approach guarantees that the hyper-rectangles contain solely points from the specific class, while also reducing their volume. We obtain a set of $50$, $100$, $200$, $250$ clusters denoted as $\mathbb{H}^{*}_{50}$$\mathbb{H}^{*}_{250}$.
Table 6 shows average volumes of the subspaces that we obtain, for two data sets. Notice consistent reduction of volume, from $\mathbb{H}_{sh}$ to $\mathbb{H}^{*}_{50}$ $\mathbb{H}^{*}_{250}$ and ultimately to $\mathbb{H}^{*}_{\epsilon=0.005}$. There are several orders of magnitude between the largest and the smallest subspace.
Hyper-rectangles Construction method Avg. volume of $\mathbb{H}$ RUAR Set cardinality of $\mathbb{H}^{*}$ RUAR Avg. volume of $\mathbb{H}$ Medical Set cardinality of $\mathbb{H}^{*}$ Medical
$\mathbb{H}_{sh}$
Hyper-rectangle shrunk to exclude all negative examples
7.55e-11
1
2.6e-09
1
$\mathbb{H}^{*}_{50}$
Set of hyper-rectangles on 50 clusters
1.02e-16
50
6.56e-15
50
$\mathbb{H}^{*}_{100}$
Set of hyper-rectangles on 100 clusters
6.23e-18
100
3.25e-17
100
$\mathbb{H}^{*}_{200}$
Set of hyper-rectangles on 200 clusters
3.31e-20
200
4.67e-19
200
$\mathbb{H}^{*}_{250}$
Set of hyper-rectangles on 250 clusters
6.42e-22
250
2.42e-20
250
$\mathbb{H}^{*}_{\epsilon=0.005}$
Set of $\epsilon$-cubes around all positive dataset sentences, $\epsilon=0.005$
1.00e-60
3400
1.00e-60
989
Table 6: Sets of geomtetric subspaces used in the experiments, their cardinality and average volumes of hyper-rectangles. All shapes are eigenspace rotated for better precision.
Verifiability Range for Subspaces.
For each set of hyper-rectangles and the given network, we pass them to the verifier (ERAN) and measure verifiability. Table 7 shows that the shrunk hyper-rectangle $\mathbb{H}_{sh}$ achieves zero verifiability, and the various clustered hyper-rectangles ($\mathbb{H}^{*}_{50}$, $\mathbb{H}^{*}_{100}$ $\mathbb{H}^{*}_{200}$, $\mathbb{H}^{*}_{250}$) achieve at most negligible verifiability. We call this effect low verifiability of the subspaces. In contrast, the baseline $\mathbb{H}^{*}_{\epsilon=0.005}$ achieves up to $99.60\%$ verifiability. This suggests that $\epsilon=0.005$ is a good benchmark for a different extreme. Table 6 can give us an intuition of why $\mathbb{H}^{*}_{\epsilon=0.005}$ has notably higher verifiability than the other hyper-rectangles: the volume of $\mathbb{H}^{*}_{\epsilon=0.005}$ is several orders of magnitude smaller.
@@
Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}^{*}_{50}}$ $\boldsymbol{\mathbb{H}^{*}_{100}}$ $\boldsymbol{\mathbb{H}^{*}_{200}}$ $\boldsymbol{\mathbb{H}^{*}_{250}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.005}}$
RUAR $N_{base}$ 0.00 0.00 1.33 0.52 0.41 0.00 88.67
Medical $N_{base}$ 0.00 0.00 0.00 2.10 4.08 5.00 97.86
@@
Table 7: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a selection of geometric subspaces; for ERAN verifier.
Verifiability-Generalisability Trade-off for Geometric Subspaces
Tables 6 and 7 suggest that smaller subspaces are more verifiable. One may also conjecture that they are less generalisable (as they will contain fewer embedded sentences). We confirm this via experiments; we are particularly interested in understanding how quickly generalisability deteriorates as verifiability increases.
To test generalisability, for each data set, we algorithmically generate a new data set $\mathcal{A}$ containing its semantic perturbations, using the method described in Section 0.3.1. It will be convenient to refer to parts of that set that correspond to perturbations of sentences belonging to positive and negative classes, respectively. In such cases, we will be using the notation $\mathcal{A}^{pos}$ and $\mathcal{A}^{neg}$. Note that in this section, we will work only with $\mathcal{A}^{pos}$. This choice is motivated by the nature of the chosen data sets: both Medical and R-U-A-Robot sentences split into:
  • a negative class that represents “all other sentences". These “other sentences" are not grouped by any specific semantic meaning and therefore do not form one coherent semantic category.
However Section 0.5 will make use of $\mathcal{A}^{neg}$ in the context of falsifiability of verified subspaces.
[…]
For generalisability, Table 8 shows that the most verifiable subspace $\mathbb{H}^{*}_{\epsilon=0.005}$ is the least generalisable. This means $\mathbb{H}^{*}_{\epsilon=0.005}$ may not contain any valid new sentences apart from the one for which it was formed! At the same time, $\mathbb{H}^{*}_{\epsilon=0.05}$ has up to $48\%$ of generalisability at the expense of only up to $5\%$ of verifiability (cf. Table 7). The effect of the generalisability vs verifiability trade-off can thus be rather severe for geometric subspaces.
This experiment demonstrates the importance of using the generalisability metric, as with only taking into account the verifiability of the subspaces one would choose $\mathbb{H}^{*}_{\epsilon=0.005}$, obtaining mathematically sound but pragmatically useless results. We argue that this is a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future.
Dataset Hyper-rectangles Avg. Volume of $\mathbb{H}$ Generalisability (%) Generalisability (# of sentences) Total Sentences in $\mathcal{A}^{pos}$
RUAR
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
1.95
2821
144500
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
38.47
55592
144500
$\mathbb{H}_{sh}$
7.55e-11
50.91
73561
144500
Medical
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
0.09
10
11209
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
28.49
3194
11209
$\mathbb{H}_{sh}$
2.6e-09
37.13
4162
11209
Table 8: Generalisability of the selected geometric subspaces $\mathbb{H}^{*}_{\epsilon=0.005}$, $\mathbb{H}^{*}_{\epsilon=0.05}$ and $\mathbb{H}_{sh}$, measured on the sets of semantic perturbations $\mathcal{A}^{pos}_{RUAR}$ and $\mathcal{A}^{pos}_{medical}$.
0.4.2 Semantic Subspaces and Generalisability-Verifiability Trade-off
The previous subsection has shown that the generalisability-verifiability trade-off is not resolvable by geometric manipulations alone. In this section we argue that using semantic subspaces can help to improve the effects of the trade-off. The main hypothesis that we are testing is: semantic subspaces permit drawing more precise shapes, and this in turn must improve both verifiability and generalisability.
Verifiability of Semantic Subspaces
We will use the construction given in Definition 3. As Table 9 illustrates, we construct several semantic hyper-rectangles on sentences of the positive class using character-level ($\mathbb{H}^{*}_{char}$, $\mathbb{H}^{*}_{del.}$, $\mathbb{H}^{*}_{ins.}$, $\mathbb{H}^{*}_{rep.}$, $\mathbb{H}^{*}_{repl.}$, $\mathbb{H}^{*}_{swap.}$), word-level ($\mathbb{H}^{*}_{word}$) and sentence-level perturbations ($\mathbb{H}^{*}_{pj}$). The subscripts ${}_{char}$ and ${}_{word}$ refer to the kind of perturbation algorithm, while ${}_{del.}$, ${}_{ins.}$, ${}_{rep.}$, ${}_{repl.}$, ${}_{swap.}$ and ${}_{pj}$ refer to the type of perturbation, where pj stands for Polyjuice (see Section 0.3.1). Notice comparable volumes of all these shapes, and compare with $\mathbb{H}^{*}_{\epsilon=0.05}$.
Hyper-rectangles Construction method Avg. volume of $\mathbb{H}$ RUAR Set cardinality of $\mathbb{H}^{*}$ RUAR Avg. volume of $\mathbb{H}$ Medical Set cardinality of $\mathbb{H}^{*}$ Medical
$\mathbb{H}^{*}_{char}$
Set of hyper-rectangles for character perturbations
1.54e-30
3400
7.66e-31
989
$\mathbb{H}^{*}_{word}$
Set of hyper-rectangles for word perturbations
1.28e-30
3400
$\mathbb{H}^{*}_{pj}$
Set of hyper-rectangles for polyjuice sentence perturbations
2.01e-28
989
$\mathbb{H}^{*}_{swap.}$
Set of hyper-rectangles for swapping perturbations
1.57e-31
3400
3.42e-31
989
$\mathbb{H}^{*}_{repl.}$
Set of hyper-rectangles for replacement perturbations
9.84e-31
3400
3.43e-31
989
$\mathbb{H}^{*}_{del.}$
Set of hyper-rectangles for deletion perturbations
3.46e-31
3400
1.24e-32
989
$\mathbb{H}^{*}_{ins.}$
Set of hyper-rectangles for insertion perturbations
3.21e-31
3400
9.11e-33
989
$\mathbb{H}^{*}_{rep.}$
Set of hyper-rectangles for repetition perturbations
1.56e-31
3400
1.06e-32
989
$\mathbb{H}^{*}_{\epsilon=0.05}$
Set of $\epsilon$-cubes generated to match the volume of the hyper-rectangles
1e-30
3400
1e-30
989
We pass each set of hyper-rectangles and the network $N_{base}$ to the verifiers ERAN and Marabou to measure verifiability of the subspaces. Table 10 illustrates the verification results obtained using ERAN. From the table, we can infer that the verifiability of our semantic hyper-rectangles is indeed higher than that of the geomtrically-defined hyper-rectangles (Table 7). Furthermore, our semantic hyper-rectangles, while unable to reach the verifiability of $\mathbb{H}^{*}_{\epsilon=0.005}$, achieve notable higher verification than its counterpart of comparable volume $\mathbb{H}^{*}_{\epsilon=0.05}$. From this experiment, we conclude that not only volume, but also precision of the subspaces has an impact on their verifiability.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{base}$ 0.00 1.80 0.87 1.62 2.63 1.66 0.94 2.07 -
Medical $N_{base}$ 5.00 - 39.71 39.62 44.66 48.71 37.49 42.60 50.09
Table 10: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for ERAN verifier.
Following these results, Table 11 reports the verification results using Marabou instead of ERAN. As shown, Marabou is able to verify up to $66.83\%$ ($\mathbb{H}^{*}_{rep.}$), while ERAN achieves at most $50.09\%$. This shows that complete and precise verification (Marabou) outperforms abstract interpretation (ERAN). Overall, the Marabou experiment confirms the trends of improved verifiability shown by ERAN and thus confirms our hypothesis about importance of shape precision.
Dataset Model $\boldsymbol{\mathbb{H}^{*}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}^{*}_{word}}$ $\boldsymbol{\mathbb{H}^{*}_{char}}$ $\boldsymbol{\mathbb{H}^{*}_{del.}}$ $\boldsymbol{\mathbb{H}^{*}_{ins.}}$ $\boldsymbol{\mathbb{H}^{*}_{rep.}}$ $\boldsymbol{\mathbb{H}^{*}_{repl.}}$ $\boldsymbol{\mathbb{H}^{*}_{swap.}}$ $\boldsymbol{\mathbb{H}^{*}_{pj}}$
RUAR $N_{base}$ 1.79 11.69 4.88 4.35 9.72 9.46 5.65 8.07 -
Medical $N_{base}$ 37.96 - 64.03 64.15 64.65 66.83 64.75 64.36 61.57
Table 11: Verifiability of the baseline DNN on the R-U-A-Robot and the Medical datasets, for a slection of semantic subspaces; for Marabou verifier.
Generalisability of Semantic Subspaces
It remains to establish whether the more verifiable semantic subspaces are also more generalisable. Table 12 augments Table 8 with results for best-verifiable semantic subspaces, $\mathbb{H}^{*}_{word}$ and $\mathbb{H}^{*}_{pj}$. It shows that these semantic subpspaces are also the most generalisable, containing, respectively, $47.67\%$ and $28.74\%$ of the unseen sentences.
[…]
Dataset Hyper-rectangles Avg. Volume of $\mathbb{H}$ Generalisability (%) Generalisability (# of sentences) Total Sentences in $\mathcal{A}^{pos}$
RUAR
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
1.95
2821
144500
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
38.47
55592
144500
$\mathbb{H}^{*}_{word}$
1.28e-30
47.67 68882
144500
Medical
$\mathbb{H}^{*}_{\epsilon=0.005}$
1.00e-60
0.09
10
11209
$\mathbb{H}^{*}_{\epsilon=0.05}$
1.00e-30
28.49
3194
11209
$\mathbb{H}^{*}_{pj}$
2.01e-28
28.74 3222
11209
Table 12: Generalisability of the selected geometric subspaces $\mathbb{H}^{*}_{\epsilon=0.005}$ and $\mathbb{H}^{*}_{\epsilon=0.05}$ and the semantic subspaces $\mathbb{H}^{*}_{word}$ and $\mathbb{H}^{*}_{pj}$, measured on the sets of semantic perturbations $\mathcal{A}^{pos}_{RUAR}$ and $\mathcal{A}^{pos}_{medical}$. Note that the generalisability of $\mathbb{H}^{*}_{sh}$ (Table 8), despite it having the volume 19 order of magnitudes bigger, is only $3\%$ greater than $\mathbb{H}^{*}_{word}$.
0.4.3 Adversarial Training on Semantic Subspaces
In this section we aim to study the effects that adversarial training methods have on the verifiability of the previously defined subspaces. By comparing a range of different training approaches, we show in this section that our new method of adversarial training based on semantic subspaces is the most efficient.
Three kinds of training are deployed in this section:

### v3

@@ 4.5 Adversarial Training on Semantic Subspaces
See Tables 16 & 19 for full listing of the networks we obtain in this way. We call DNNs of second and third type robustly trained networks. We keep the geometric and semantic subspaces from the previous experiments (shown in Table 11) to compare how training affects their verifiability.
Following the same evaluation methodology of experiments as in Sections 4.2 and 4.4.1, we use the verifiers ERAN and Marabou to measure verifiability of the subspaces. Table 16 reports accuracy of the robustly trained networks, while the verification results are presented in Tables 17 and 18. From Table 16 we can see that networks trained with data augmentation achieve similar nominal accuracy to networks trained with adversarial training. However, the most prominent difference is exposed in Tables 17 and 18: adversarial training effectively improves the verifiability of the networks, while data augmentation actually decreases it.
Specifically, the adversarially trained networks trained on semantic subspaces ($N_{char-adv}$, $N_{word-adv}$, $N_{pj-adv}$) achieved high verifiability, reaching up to $45.87\%$ for RUAR and up to $83.48\%$ for the Medical dataset. This constitutes a significant improvement of the verifiability results compared to $N_{base}$. Looking at nuances, there does not seem to be a single winner subspace when it comes to adversarial training, and indeed in some cases $\mathbb{H}_{\epsilon=0.05}$ wins over more precise subspaces. All of the subspaces in Table 11 have very similar volume, which accounts for improved performance across all experiments. The particular peaks in performance then come down to particularities of a specific semantic attack that was used while training. For example, the best performing networks are those trained with Polyjuice attack, the strongest form of attack in our range. Thus, if the kind of attack is known in advance, the precision of hyper-rectangles can be further tuned.
Model Dataset Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
$N_{char-aug}$ $\mathcal{Y}\cup\mathcal{A}^{5}_{char}(\mathcal{Y}^{pos})$ 95.62 ± 0.26% 93.20 ± 0.35% 99.08 ± 0.06% 93.46 ± 0.30%
$N_{word-aug}$ $\mathcal{Y}\cup\mathcal{A}^{6}_{word}(\mathcal{Y}^{pos})$ 98.57 ± 0.06% 94.59 ± 0.36% - -
$N_{pj-aug}$ $\mathcal{Y}\cup\mathcal{A}^{5}_{pj}(\mathcal{Y}^{pos})$ 98.19 ± 0.09% 93.19 ± 0.39%
$N_{char-adv}$ $\mathcal{Y}$ 93.26 ± 0.19% 92.51 ± 0.38% 96.27 ± 0.05% 95.09 ± 0.16%
$N_{word-adv}$ $\mathcal{Y}$ 93.68 ± 0.16% 92.37 ± 0.29% - -
$N_{pj-adv}$ $\mathcal{Y}$ 95.05 ± 0.19% 93.49 ± 0.32%
$N_{\epsilon=0.05-adv}$ $\mathcal{Y}$ 94.01 ± 0.17% 92.24 ± 0.19% 96.05 ± 0.09% 95.04 ± 0.24%
Table 16: Accuracy of the robustly trained DNNs on the RUAR and the Medical datasets. $\mathcal{Y}$ stands for either RUAR or Medical depending on the column.
@@
Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
RUAR $N_{char-aug}$ 0.00% 0.24% 0.00% 0.51% 1.38% 1.09% 0.35% 1.06% -
$N_{word-aug}$ 0.00% 0.24% 0.00% 0.42% 0.31% 0.57% 0.25% 0.92% -
$N_{char-adv}$ 0.00% 8.97% 4.43% 4.81% 9.86% 11.3% 6.91% 8.51% -
$N_{word-adv}$ 0.04% 10.75% 4.05% 4.36% 8.60% 9.52% 6.81% 7.45% -
$N_{\epsilon=0.05-adv}$ 0.12% 10.16% 4.18% 4.04% 8.91% 10.17% 6.52% 7.36% -
Medical $N_{char-aug}$ 0.00% - 7.59% 5.28% 12.84% 11.05% 7.92% 7.40% 26.97%
$N_{pj-aug}$ 0.00% - 10.31% 8.49% 15.67% 14.90% 9.18% 10.58% 28.59%
$N_{char-adv}$ 5.28% - 50.12% 49.78% 53.99% 57.76% 48.02% 52.07% 55.44%
$N_{pj-adv}$ 2.83% - 47.11% 46.14% 52.12% 56.14% 44.59% 48.27% 57.36%
$N_{\epsilon=0.05-adv}$ 8.68% - 51.60% 50.31% 55.67% 58.52% 50.10% 53.65% 59.76%
@@
Table 17: Verifiability of the robustly trained DNNs on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the ERAN verifier.
Dataset Model $\boldsymbol{\mathbb{H}_{\epsilon=0.05}}$ $\boldsymbol{\mathbb{H}_{word}}$ $\boldsymbol{\mathbb{H}_{char}}$ $\boldsymbol{\mathbb{H}_{del.}}$ $\boldsymbol{\mathbb{H}_{ins.}}$ $\boldsymbol{\mathbb{H}_{rep.}}$ $\boldsymbol{\mathbb{H}_{repl.}}$ $\boldsymbol{\mathbb{H}_{swap.}}$ $\boldsymbol{\mathbb{H}_{pj}}$
RUAR $N_{char-aug}$ 0.72% 13.90% 8.49% 7.92% 13.67% 15.50% 9.56% 11.88% -
$N_{word-aug}$ 0.24% 11.30% 3.87% 4.05% 8.27% 8.84% 5.71% 7.72% -
$N_{char-adv}$ 7.37% 41.93% 30.41% 30.23% 38.20% 45.87% 32.74% 36.62% -
$N_{word-adv}$ 12.17% 45.12% 25.82% 25.39% 33.85% 37.45% 26.87% 30.99% -
$N_{\epsilon=0.05-adv}$ 18.46% 41.93% 21.99% 20.32% 28.13% 32.83% 23.52% 26.74% -
Medical $N_{char-aug}$ 1.14% - 37.05% 35.29% 41.50% 42.47% 34.89% 37.94% 49.65%
$N_{pj-aug}$ 5.77% - 39.00% 38.66% 42.28% 44.22% 37.29% 39.03% 38.22%
$N_{char-adv}$ 51.70% - 77.59% 77.25% 77.50% 77.98% 77.92% 78.67% 76.58%
$N_{pj-adv}$ 57.45% - 81.94% 81.47% 82.31% 83.48% 82.47% 82.72% 82.24%
$N_{\epsilon=0.05-adv}$ 62.57% - 79.32% 78.57% 78.70% 80.21% 79.40% 80.76% 66.22%
Table 18: Verifiability of the DNNs trained for robustness on the RUAR and the Medical datasets, for a selection of semantic subspaces; using the Marabou verifier.
Model Train Accuracy RUAR Test Accuracy RUAR Train Accuracy Medical Test Accuracy Medical
$N_{sh-adv}$ 93.39 ± 0.22% 92.96 ± 0.13% 96.14 ± 0.12% 94.29 ± 0.26%
$N_{50-adv}$ 94.32 ± 0.14% 93.49 ± 0.19% 95.56 ± 0.20% 95.15 ± 0.12%
$N_{100-adv}$ 94.88 ± 0.04% 94.18 ± 0.24% 95.71 ± 0.11% 95.47 ± 0.16%
$N_{200-adv}$ 95.09 ± 0.09% 94.45 ± 0.14% 95.85 ± 0.05% 95.43 ± 0.10%
$N_{250-adv}$ 95.22 ± 0.08% 94.22 ± 0.23% 96.07 ± 0.13% 95.38 ± 0.22%
$N_{\epsilon=0.005-adv}$ 93.48 ± 0.21% 91.59 ± 0.07% 96.24 ± 0.04% 95.13 ± 0.09%
Table 19: Accuracy of the DNNs trained adversarially on the RUAR and the Medical datasets.
Dataset Model $\boldsymbol{\mathbb{H}_{sh}}$ $\boldsymbol{\mathbb{H}_{50}}$ $\boldsymbol{\mathbb{H}_{100}}$ $\boldsymbol{\mathbb{H}_{200}}$ $\boldsymbol{\mathbb{H}_{250}}$ $\boldsymbol{\mathbb{H}_{\epsilon=0.005}}$
RUAR $N_{sh-adv}$ 0.00% 0.00% 1.33% 0.52% 0.41% 88.62%
$N_{50-adv}$ 0.00% 0.00% 0.00% 0.00% 0.41% 90.02%
$N_{100-adv}$ 0.00% 0.00% 0.00% 0.00% 0.41% 92.74%
$N_{200-adv}$ 0.00% 0.00% 0.00% 0.00% 0.08% 93.54%
$N_{250-adv}$ 0.00% 0.00% 0.00% 0.00% 0.33% 93.86%
$N_{\epsilon=0.005-adv}$ 0.00% 0.00% 0.00% 0.00% 0.33% 98.22%
Medical $N_{sh-adv}$ 0.00% 0.00% 0.00% 2.50% 4.40% 97.47%
$N_{50-adv}$ 0.00% 0.00% 1.08% 3.60% 6.00% 98.79%
$N_{100-adv}$ 0.00% 0.00% 1.08% 3.00% 5.04% 99.09%
$N_{200-adv}$ 0.00% 0.00% 1.08% 2.90% 4.96% 99.05%
$N_{250-adv}$ 0.00% 0.00% 0.00% 2.90% 4.40% 98.73%
$N_{\epsilon=0.005-adv}$ 0.00% 0.00% 0.00% 2.30% 4.32% 99.60%
Table 20: Verifiability of the DNNs trained adversarially on the RUAR and the Medical datasets, for a selection of geometric subspaces; using the ERAN verifier.
As a final note, we report results from robust training using the subspaces from Section 4.2 in Table 8. Table 19 reports the accuracy and the details of the robustly trained networks on those subspaces, while the verification results are presented in Table 20. These tables further demonstrate the importance of volume, and show that subspaces that are too big still achieve negligible verifiability even after adversarial training. Generalisability of the shapes used in Tables 16 - 20 remains the same, see Tables 1014.
4.5.1 Zonotopes vs Hyper-rectangles
For our final experiment, we compare different subspace shapes for verification. Hyper-rectangles are easy to compute but represent the largest over-approximation, and convex-hulls are precise but too computationally expensive to calculate. Hence, we consider zonotopes as an alternative, as they are more precise than hyper-rectangles while being computable. Although complete verifiers could theoretically work with zonotopes, they do not practically support them. Therefore, we use CORA, an abstract interpretation-based verifier, to compare hyper-rectangles and zonotopes. Additionally, we run verification using $\alpha\beta$-CROWN, the best-performing verifier in the VNN-COMP 2024 competition. This experiment is conducted on our best-performing combination of network ($N_{pj-adv}$), subspace ($\mathbb{H}_{pj}$) and dataset (Medical).
Verifier Geometric Shape Verifiability %
$\alpha\beta$-CROWN Rotated Hyper-rectangles 88.41
Marabou Rotated Hyper-rectangles 82.24
ERAN Rotated Hyper-rectangles 57.36
CORA Zonotopes 55.73
CORA Rotated Hyper-rectangles 29.10
Table 21: Comparison of different subspace shapes for verification and verifiers for $N_{pj-adv}$ and $\mathbb{H}_{pj}$ on the Medical dataset.
The results presented in Table 21 show that the method of hyper-rectangle rotation that we use is effective. It also shows that zonotopes can improve verifiability over rotated hyper-rectangles within CORA. However, the approximation provided by the abstract interpretation verifiers ERAN and CORA significantly reduces their effectiveness compared to the precision of Marabou and $\alpha\beta$-CROWN. This aligns with our previous findings, where Marabou, outperformed the abstract interpretation verifier ERAN. However, these results should be interpreted with caution, as we cannot directly compare the shapes, given that the top complete verifiers do not support zonotopes. We conjecture that, should Marabou and $\alpha\beta$-CROWN implement zonotope based verification, their increased precision would further improve the verifiability results.
5 NLP Case Studies
The purpose of this section is two-fold. Firstly, the case studies we present here apply the NLP Verification Pipeline set out in Section 2.6 using a wider range of NLP tools. Notably, in this section we try different LLMs to embed sentences and replace Polyjuice with the LLM vicuna-13b [^8] Using the following API: https://replicate.com/replicate/vicuna-13b/api., a state-of-the-art open source chatbot trained by fine-tuning LLaMA [138] on user-shared conversations collected from ShareGPT  [^9] https://sharegpt.com/. For further details, please refer to Section 3.1. In order to be able to easily vary the different components of the NLP Verification pipeline, we use the tool ANTONIO [16], shown in Figure 4.
Figure 4: Tool ANTONIO that implements a modular approach to the NLP verification pipeline used in this paper.
Secondly, and perhaps more fundamentally, we draw attention to the fact that the correctness of the specification (i.e. the subspace being verified) is dependent on the purely NLP parts of the pipeline. In particular, the parts that generate, perturb, and embed sentences. Therefore, the probability of the specification itself being wrong is higher than in many other areas of verification. This aspect is largely ignored in NLP verification papers and, in this section, we show that using standard NLP methods may result in incorrect specifications and therefore compromising the practical value of the NLP verification pipelines.
Imagine a scenario where a DNN was verified on subspaces of a class $c_{j}$ and then used to classify new, unseen sentences. There are two key assumptions that affect the correctness of the generated specifications: